A retired woman has $50,000 to invest but needs to make $6,000 a year from the interest to meet certain living expenses. One bond investment pays 15% annual interest. The rest of it she wants to put in a CD that pays 7%. Set up and solve the equation for how much the woman should invest in each option to sustain a $6,000 annual return.
Answer by addingup(3677)

Let the amount invested at 15% be x:
0.15x + 0.07(50,000-x) = 6,000
0.15x + 3,500 - 0.07x = 6,000
0.08x = 2,500
x = 31,250 this is the amount invested at 15%
Thus, the amount invested at 7% is 50,000 - 31,250 = 18,750
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
0.15 * 31,250 = 4,687.50
0.07 * 18,750 = 1,312.50
. . . . . . . .-----------
Total .. . . . $6,000 Correct